On 12-31-15, Acme purchased a machine. Acme signed a $800,000 zero-interest bearing note. The note is payable in full on 12-31-17. Assume an acceptable interest rate on similar notes was 4%. On 12-31-15, Acme incurred and paid $12,000 to have the machine installed in its sales office. In this problem, you can ignore depreciation – we’ll get to that in chapter 11. Prepare the entries Acme should make related to this note on:

  1. 12-31-15.
  2. 12-31-16.
  3. 12-31-17.
Answer #1

Zero interest notes are issued at a lower price than their face value / future value. Here face value / future value is $ 800,000

Acme has purchased machine against the zero interest note whose face value is $ 800,000. To find out the cost of machine we need to find out present value of zero interest bond.

Present Value (PV) = FV / (1 + i%)n

where FV = future value, i% = the interest rate and n= the term in years

Here in current scenario:

FV = $ 800,000

i% = 4% ( as interest on similar bond is 4%)

n= 2 years

PV= $ 800,000 /(1 + 4%)2 = $ 739,645

Cost of machine will be recorded at $ 739,645 plus installation charges of $ 12,000

Difference between PV and FV is the interest / discount on bonds and will be amortized over 2 years.

Discount / Interest on bond = FV - PV = $ 800,000 - 739,645 = $ 60,355

Amortization of interest expense per year = (FV - PV) / No. of years for bond payment

Amortization of Interest expense per year = ($ 800,000 - 739,645) / 2 = $ 30,177.5

Accounting entries that Acme should make are as under:

Date General Journal Debit Credit
31-Dec-15 Machine A/c $           751,645
Discount on notes payable A/c $             60,355
To Notes Payable A/c $                     800,000
To Cash A/c $                        12,000
(being machine purchased against zero interest note payable on 12-31-17 FV $ 800,000, installation charges paid in cash $ 12,000)
31-Dec-16 Interest Expense A/c $       30,177.50
To Discount on notes payable A/c $                  30,177.50
(being interest amount for the year on bond charged to expense)
31-Dec-17 Interest Expense A/c $       30,177.50
To Discount on notes payable A/c $                  30,177.50
(being interest amount for the year on bond charged to expense)
31-Dec-17 Notes Payable A/c $           800,000
To Cash A/c $                     800,000
(being cash paid against bond on due date)

Explanation:

Installation charges will be debited to cost of machinery as all the expenses paid to put the machine to use should form part of cost of machine

Discount on notes payable account is a balance sheet contra liability account, it will be netted off against the notes payable account to display the net liability.